OpenStudy (anonymous):

Which of these ordered pairs is a solution to the linear inequality y – 4x ≤ –6? A. (–2, 4) B. (1, –2) C. (1, 3) D. (2, 3) I also need help in Linear Inequalities/Equations as well. All of it.

OpenStudy (anonymous):

B.

OpenStudy (texaschic101):

y - 4x <= -6 ....(1,-2) -2 - 4(1) <= -6 -2 - 4 <= -6 -6 <= -6 (correct) AnswerMyQuestions is correct...answer is B

OpenStudy (anonymous):

Thanks for the medal! :D

OpenStudy (texaschic101):

you deserved it :)

OpenStudy (anonymous):

Thank you! :D Really helped!

OpenStudy (texaschic101):

glad to help :)